OpenStudy (peterg2001):

Which of the following will form a right triangle? 7, 24, 25 7, 23, 24 7, 25, 26

OpenStudy (peterg2001):

@xX_GamerGirl_Xx @S117 @EclipsedStar

OpenStudy (peterg2001):

@jjfreerider2316 @Directrix

OpenStudy (xx_gamergirl_xx):

look in msg (message)

OpenStudy (texaschic101):

a^2 + b^2 = c^2 ---- c being the biggest number 7,24,25 7^2 + 24^2 = 25^2 49 + 576 = 625 625 = 625 (bingo)....this is a right triangle